# Power-law banded random matrix ensemble as a model for quantum many-body Hamiltonians

### v2

@@ VII Concluding discussion
In this work we explored interpretations of the PLBRM ensemble as Hamiltonians of one-dimensional quantum many-body systems, either disordered or chaotic. This was motivated by the observation that this class of random matrices, in the weakly ergodic phase $\frac{1}{2}<\alpha<1$, displays the distinction between edge and bulk eigenstates that is typical of many-body quantum systems. In many-body systems with local interactions, this edge-bulk distinction leads to the eigenstate entanglement entropy having a characteristic rainbow shape, which is absent in the GOE and GUE ensembles but present in the weakly ergodic phase of the PLBRM ensemble.
@@
We proposed and compared three different ways to assign many-body product basis states to the basis vectors of the random matrices. Our Gray code labeling scheme improves on the previously used binary labeling scheme. We explained the need to apply ‘site randomization’ in order to ensure that the resulting many-body Hamiltonians are spatially uniform. Using this setup, we focused on a particular physical property of the resulting Hamiltonians, namely, the half-system bipartite eigenstate entanglement entropy $S_{\text{ent}}$. We examined the behavior of $S_{\text{ent}}$ in each of the phases. For all the labeling schemes, the common features are: (1) In the fully ergodic phase, $\alpha<\frac{1}{2}$, the eigenstate entanglement entropy saturates the bound for fully ergodic states ($S_{\text{ent}}\to S_{\text{Page}}$) over the complete spectrum . (2) In the weakly ergodic phase, $\frac{1}{2}<\alpha<1$, there is significant variation across the spectrum: the bulk and edge eigenstates have different scaling behaviors. (3) In the localized phase $\alpha>1$, the complete spectrum shows area law or at least sub-volume law behavior.
@@
For the thermodynamic limit, these behaviors imply, firstly, that a sharp change should be seen in the edge $S_{\text{ent}}$ at the boundary between the fully and weakly ergodic phases (at $\alpha=\frac{1}{2}$), and secondly, that a sharp change should be seen in the bulk $S_{\text{ent}}$ at the boundary between the weakly ergodic and localized phases (at $\alpha=1$). We have confirmed this expectation by analyzing the finite-size crossings in the $S_{\text{ent}}$ versus $\alpha$ curves.
@@
The energy-dependence is arguably the most interesting in the intermediate weakly ergodic phase. This is also the regime that models the rainbow behavior of $S_{\text{ent}}$, which was a primary motivation behind this study. We therefore performed a more detailed study of this regime. We have managed to demarcate the eigenstates into three types according to their entanglement scaling behaviors. The bulk eigenstates show entanglement entropy equal to these for maximally ergodic states ($S_{\text{ent}}\to S_{\text{Page}}$), the edge eigenstates show area law, and there is an intermediate set of eigenstates that have volume law but a deviation $S_{\text{Page}}-S_{\text{ent}}$ that is non-decreasing with system size. The structure is richer than a single demarcation between the area law eigenstates and the $S_{\text{ent}}{\to}S_{\text{ent}}$ eigenstates.
@@
Of the two exponents marking the two boundaries, $\delta_{1}$, which separates decreasing from increasing deviations from $S_{\text{Page}}$, follows the simple behavior $\delta_{1}=\alpha$. The other exponent, $\delta_{2}$, marking the area law eigenstates at the very edges of the spectrum, was numerically more challenging to calculate: our procedure only provides a lower bound. The data is consistent with $\delta_{2}$ increasing from 00 to $1$ as $\alpha$ increases from $\frac{1}{2}$ to $1$, although this behavior near $\alpha=1$ is challenging to confirm (Fig. 11). We pointed out that the result $\delta_{1}=\alpha$ has the following curious consequence. As one approaches the fully ergodic phase ($\alpha\to\frac{1}{2}+0^{+}$), one might naively expect that the number of eigenstates that deviate from $S_{\text{Page}}$ should vanish. However, our result $\delta_{1}=\alpha$ indicates that $\sim N^{1/2}$ eigenstates show non-vanishing deviation from the Page value, for $\alpha\to\frac{1}{2}+0^{+}$. This is a vanishing fraction, but a diverging number. This type of transition is similar to the quantum Zeno transition for the inverse participation ratio, $\text{IPR}=c_{2}N^{-D_{2}}$, where the fractal dimension $D_{2}$, the analogue of our critical exponent $\delta_{1}$, does not undergo any transition, while the prefactor $c_{2}$ (the analogue of our $S_{\text{Page}}-S_{\text{ent}}$) develops a jump at the transition from a finite to zero value [114].
@@
(1) Many-body systems which thermalize have similar entanglement profile as our weakly ergodic phase. It would be interesting to try to demarcate area-law edge states from volume-law bulk states in such a case; we are not aware of a many-body system where such a boundary has been quantitatively identified. An additional interesting question for many-body systems is: whether there might be an intermediate set of eigenstates with non-vanishing deviation as in the present case, or other type of intermediate behavior.
@@
(5) Another interesting question of convention is the issue of labeling schemes connecting PLBRM indices to many-body basis states. For our analysis of the energy-dependence of $S_{\text{ent}}$ scaling in the weakly ergodic phase (Sec. VI), we focused on the Gray code scheme. It is possible that these results might depend on the labeling scheme, e.g., the random or binary schemes might have different demarcation behaviors between edge and bulk, or different values of $\delta_{1}$ and $\delta_{2}$. It would therefore be fruitful in future work to map out the dependence on labeling scheme.
@@ Acknowledgements.
MH acknowledges support from the Deutsche Forschungsgemeinschaft under grant SFB 1143 (project-id 247310070). IMK acknowledges support by the European Research Council under the European Union’s Seventh Framework Program Synergy ERC-2018-SyG HERO-810451.

### v3

@@ VII Concluding discussion
In this work we explored interpretations of the PLBRM ensemble as Hamiltonians of one-dimensional quantum many-body systems, either disordered or chaotic. This was motivated by the observation that this class of random matrices, in the weakly ergodic phase $\frac{1}{2}<\alpha<1$, displays the distinction between edge and bulk eigenstates that is typical of many-body quantum systems. In many-body systems with local interactions, this spectral edge-bulk distinction leads to the eigenstate entanglement entropy having a characteristic rainbow shape, which is absent in the GOE and GUE ensembles but present in the weakly ergodic phase of the PLBRM ensemble.
@@
We proposed and compared three different ways to assign many-body product basis states to the basis vectors of the random matrices. Our Gray code labeling scheme improves on the previously used binary labeling scheme. We explained the need to apply “site randomization” to ensure that the resulting many-body Hamiltonians are spatially uniform. Using this setup, we focused on a particular physical property of the resulting Hamiltonians, namely, the half-system bipartite eigenstate entanglement entropy $S_{\text{ent}}$. We examined the behavior of $S_{\text{ent}}$ in each of the phases. For all the labeling schemes, the common features are: (1) In the fully ergodic phase, $\alpha<\frac{1}{2}$, the eigenstate entanglement entropy saturates the bound for fully ergodic states ($S_{\text{ent}}\to S_{\text{Page}}$) over the complete spectrum. (2) In the weakly ergodic phase, $\frac{1}{2}<\alpha<1$, there is significant variation across the spectrum: the spectral bulk and spectral edge eigenstates have different scaling behaviors. (3) In the localized phase $\alpha>1$, the complete spectrum shows area law or at least sub-volume law behavior.
@@
For the thermodynamic limit, these behaviors imply, firstly, that a sharp change should be seen in the spectral edge $S_{\text{ent}}$ at the critical point between the fully and weakly ergodic phases (at $\alpha=\frac{1}{2}$), and secondly, that a sharp change should be seen in the bulk $S_{\text{ent}}$ at the critical point between the weakly ergodic and localized phases (at $\alpha=1$). We have confirmed this expectation by analyzing the finite-size crossings in the $S_{\text{ent}}$ versus $\alpha$ curves.
@@
The energy-dependence is arguably the most interesting in the intermediate weakly ergodic phase. This is also the regime that models the rainbow behavior of $S_{\text{ent}}$, which was a primary motivation behind this study. We therefore performed a more detailed study of this regime. We have managed to demarcate the eigenstates into three types according to their entanglement scaling behaviors. The spectral bulk eigenstates show entanglement entropy equal to these for maximally ergodic states ($S_{\text{ent}}\to S_{\text{Page}}$), the spectral edge eigenstates show area law, and there is an intermediate set of eigenstates that have volume law but a deviation $S_{\text{Page}}-S_{\text{ent}}$ that is non-decreasing with system size. The structure is richer than a single demarcation between the area law eigenstates and the $S_{\text{ent}}{\to}S_{\text{ent}}$ eigenstates.
@@
Of the two exponents marking the two boundaries in the spectrum, $\delta_{1}$, which separates decreasing from increasing deviations from $S_{\text{Page}}$, follows the simple behavior $\delta_{1}=\alpha$. The other exponent, $\delta_{2}$, marking the area law eigenstates at the very edges of the spectrum, was numerically more challenging to calculate: our procedure only provides a lower bound. The data is consistent with $\delta_{2}$ increasing from 0 to $1$ as $\alpha$ increases from $\frac{1}{2}$ to $1$, although this behavior near $\alpha=1$ is challenging to confirm (Fig. 11). We pointed out that the result $\delta_{1}=\alpha$ has the following curious consequence. As one approaches the fully ergodic phase ($\alpha\to\frac{1}{2}+0^{+}$), one might naively expect that the number of eigenstates that deviate from $S_{\text{Page}}$ should vanish. However, our result $\delta_{1}=\alpha$ indicates that $\sim N^{1/2}$ eigenstates show nonvanishing deviation from the Page value, for $\alpha\to\frac{1}{2}+0^{+}$. This is a vanishing fraction, but a diverging number. This type of transition is similar to the quantum Zeno transition for the inverse participation ratio, $\text{IPR}=c_{2}N^{-D_{2}}$, where the fractal dimension $D_{2}$, the analog of our critical exponent $\delta_{1}$, does not undergo any transition, while the prefactor $c_{2}$ (the analog of our $S_{\text{Page}}-S_{\text{ent}}$) develops a jump at the transition from a finite to zero value [115].
Recent analytic work of interest to the present line of inquiry includes studies of entanglement in Gaussian eigenstates [116, 67, 117, 118, 96], and a treatment of the crossover between edge and bulk behavior in conformal field theory and free-fermion systems [119]. We are not aware of cases where the spectral boundaries between different eigenstate entanglement behaviors have been quantified.
@@
(1) Many-body systems which thermalize have similar entanglement profiles as our weakly ergodic phase. It would be interesting to try to demarcate area-law edge states from volume-law bulk states in such a case; we are not aware of a many-body system where such a boundary has been quantitatively identified. An additional interesting question for many-body systems is: whether there might be an intermediate set of eigenstates with nonvanishing deviation as in the present case, or other type of intermediate behavior.
@@
(5) Another interesting question of convention is the issue of labeling schemes connecting PLBRM indices to many-body basis states. For our analysis of the energy-dependence of $S_{\text{ent}}$ scaling in the weakly ergodic phase (Sec. VI), we focused on the Gray code scheme. It is possible that these results might depend on the labeling scheme, e.g., the random or binary schemes might have different demarcation behaviors between spectral edge and bulk, or different values of $\delta_{1}$ and $\delta_{2}$. It would therefore be fruitful in future work to map out the dependence on labeling scheme.
@@
(7) In this work we focused on the PLBRM class as a minimal model for displaying and analyzing the entanglement rainbow structure. One can extend the GOE/GUE ensembles in various other ways to incorporate features of many-body Hamiltonians, e.g., one could include sparsity, locality etc.
@@ Acknowledgements.
MH acknowledges support from the Deutsche Forschungsgemeinschaft under Grant No. SFB 1143 (Project ID No 247310070). IMK acknowledges support by the European Research Council under the European Union’s Seventh Framework Program Synergy ERC-2018-SyG HERO-810451. MH thanks M. Rigol and L. Vidmar for fruitful discussions and feedback.